OpenStudy (anonymous):

the result of a number,when decreased by 20% ,is 192.find this number

OpenStudy (anonymous):

@shamim @Smita12 @SchoolHelpPlease169 @ganeshie8 @queelius @iambatman @ikram002p @inowalst

OpenStudy (inowalst):

@iambatman ??

OpenStudy (anonymous):

can help/?

OpenStudy (shamim):

Let the number is x

OpenStudy (aaronandyson):

Let the number be "x": Equation: x - 0.20x = 192?

OpenStudy (shamim):

So amount of decrease=x*20%=?

OpenStudy (shamim):

Response plz!!!

OpenStudy (shamim):

Anyway x*20%=x*20/100=0.2x Right?

OpenStudy (shamim):

So x-0.2x=192

OpenStudy (shamim):

x=?

OpenStudy (anonymous):

oh sorry ...wait please

OpenStudy (shamim):

Ok....

OpenStudy (anonymous):

so answer?

OpenStudy (anonymous):

answer would be what?@shamim @AaronAndyson

OpenStudy (anonymous):

@shamim

OpenStudy (anonymous):

x - 0.20x = 192

OpenStudy (anonymous):

next @Smita12

OpenStudy (anonymous):

x - \[\frac{ 1}{ 5x}\] = 192

OpenStudy (anonymous):

these are in the same line

OpenStudy (anonymous):

it should be 1x / 5

OpenStudy (anonymous):

\[\frac{ 5x - x }{ 5}\]

OpenStudy (anonymous):

its simple wait u have 20x/100

OpenStudy (anonymous):

so we canceling the "0" andcanceling 10 with 2

OpenStudy (shamim):

U r wrongv@Smita12

OpenStudy (anonymous):

no my method is diff

OpenStudy (anonymous):

so we get 0.2... next @shamim tell me from dwhere u left plzzz

OpenStudy (shamim):

x-.2x=192 x(1-.2)=192 x*.8=192 x=192/.8 x=?

OpenStudy (anonymous):

how you get 1 and .8 @shamim

OpenStudy (anonymous):

@Smita12 please countinue your procedure

OpenStudy (anonymous):

lol

OpenStudy (shamim):

U know x=x*1 Right?

OpenStudy (anonymous):

just tell me first how much did u understand

OpenStudy (anonymous):

x-0.2=192 still

OpenStudy (anonymous):

no @shamim

OpenStudy (anonymous):

it should be x - 0.2x = 192

OpenStudy (anonymous):

take x as 1

OpenStudy (shamim):

x-0.2x=192 x*1-.2x=192 x(1-.2)=192 x*.8=192 Right?

OpenStudy (anonymous):

1 - 0.2 = 0.8

OpenStudy (anonymous):

so i suppose it is noww 0.8x

OpenStudy (anonymous):

0.8x = 192

OpenStudy (anonymous):

x = 192/0.8

OpenStudy (anonymous):

thank you very much all of you

OpenStudy (leonardo0430):

^